Bellawella:

Diego is packing 280 CDs into boxes. Each box holds 8 CDs. Diego has already packed 200 CDs. Which equation can be used to find the number of boxes, B, that Diego needs in order to finish packing CDs? text him the answer Jjjones32

Imagine:

\[200+8x=280\]

Imagine:

Minus 200 from 280.

Imagine:

\[8x=80\]

Imagine:

Figure it out from there.

supie:

Since the equation says there's 280 in 1 box then says each box holds 8 so we don't know how much boxes so it would be \[280+8x=?\] So we have to find that Then the equation says "Diego has already packed 200 CDs" so since he has already did it we have to subtract 200 from 280 to get the value of 8x \[280-200=80\] So \[8x=80\] (so x=10) And yeah, like imagine said, you can figure it out from here :)

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie Since the equation says there's 280 in 1 box then says each box holds 8 so we don't know how much boxes so it would be \[280+8x=?\] So we have to find that Then the equation says "Diego has already packed 200 CDs" so since he has already did it we have to subtract 200 from 280 to get the value of 8x \[280-200=80\] So \[8x=80\] (so x=10) And yeah, like imagine said, you can figure it out from here :) \(\color{#0cbb34}{\text{End of Quote}}\) look please on the @Imagine calcule ... than 8x = 80 so how you ve wrote above 280 +80 not equal 280